One-half of Mr. Staines' age four years from now minus three-fourths of Mr. Staines' age four years ago is zero. What is Mr. Staines' present age?
Answer by ptaylor(2198)

Let x=Mr. Staine's present age
x+4=his age 4 years from now and
x-4=his age 4 years ago
Now we are told that:

(1/2)(x+4)-(3/4)(x-4)=0 simplify
(x/2)+2-(3x/4)+3=0 multiply each term by 4
2x+8-3x+12=0 or
-x+20=0 subtract 20 from each side
-x=-20 divide both sides by -1
x=20-------------------------------------Mr. Staine's present age
CK
(1/2)(20+4)-(3/4)(20-4)=0
12-12=0
0=0

Hope this helps---ptaylor